What will be the inverse of the following functions from R to R?

(a)

f: R—>R defined by f(x) = x

(b)

f: R—>R defined by f(x) = x + 1

(c)

f: R—>R defined by f(x) = – 3x+4

(d)

f: R—>R defined by f(x) = x^3

(e)

f: R—>R defined by f(x) = sin (x)


1
Expert's answer
2021-07-19T10:14:42-0400

Solution:

(a)

"f(x)=y=x"

Interchange x and y,

"x=y"

Now solving for y, we get,

"\therefore f^{-1}(x)=x"

(b)

"f(x)=y = x + 1"

Interchange x and y,

"x=y+1"

Now solving for y, we get,

"y=x-1
\\\therefore f^{-1}(x)=x-1"

(c)

"f(x)=y = \u2013 3x+4"

Interchange x and y,

"x=-3y+4"

Now solving for y, we get,

"y=\dfrac{x+4}3
\\\therefore f^{-1}(x)=\dfrac{x+4}3"

(d)

"f(x)=y = x^3"

Interchange x and y,

"x=y^3"

Now solving for y, we get,

"y=x^{1/3}
\\\therefore f^{-1}(x)=x^{1/3}"

(e)

"f(x)=y = \sin x"

Interchange x and y,

"x=\sin y"

Now solving for y, we get,

"y=\sin^{-1}x
\\\therefore f^{-1}(x)=\sin^{-1}x"